Consider the data set shown below. Find the 95% confidence interval for the slope of the regression line.

y 0 3 2 3 8 10 11

x -2 0 2 4 6 8 10

9) A) 0.94643 ± 0.27603 B) 0.94643 ± 0.33306 C) 0.94643 ± 0.28377 D) 0.94643 ± 0.36203
Answer #1

Answer:-

Given That:-

Consider the data set shown below. Find the 95% confidence interval for the slope of the regression line.

y 0 3 2 3 8 10 11

x -2 0 2 4 6 8 10

Given,

X_i Y_i x Y_i^2 X_iY_i \hat{Y_i} e_i=Y_i-\hat{Y_i}
-2 0 4 0 0 -0.3927 at x = -2 0.3927
0 3 0 9 0 1.5001 at x = 0 1.4999
2 2 4 4 4 3.3929 at x = 2 -1.3929
4 3 16 9 12 5.2857 at x = 4 -2.2857
6 8 36 64 49 7.1785 at x = 6 0.8215
8 10 64 100 80 9.0713 at x = 8 0.9287
10 11 100 121 110 10.9641 at x =10 0.0359

\bar{X}=\frac{\Sigma X_i}{n}

= 28/7

= 4

\bar{Y}=\frac{\Sigma Y_i}{n}

= 37/7

= 5.2857

\Sigma X_iY_i=254

\Sigma X_i^2=224

\hat{\beta}=\frac{\Sigma X_iY_i-n\bar{X}\bar{Y}}{\Sigma X^2_i-n\bar{X}^2}

\hat{\beta}=\frac{254-7*4*5.2857}{224-7*(4)^2}

= 106.0004/112

= 0.9464

\hat{\beta}_0=\bar{Y}-\hat{\beta}_1\bar{X}

= 5.2857 - 0.9464 * 4

= 1.5001

Simple linear regression model

Y=\beta_0+\beta_1 X+\varepsilon \sim N(0, \sigma^2)

\hat{Y}=\beta_0+\beta_1 X+\varepsilon \sim N(0, \sigma^2)

\hat{Y}=1.5001+0.9464X_i

\Sigma e^2_i=(0.3927)^2+(1.4999)^2+(-1.3929)^2+(-2.2857)^2+(0.8215)^2+(0.9287)^2+(0.0359)^2

\Sigma e^2_i=11.1072=SS_{Residual}

MS_{Residual}=\frac{SS_{Residual}}{n-2}

= 11.1072/7-2

= 2.2213

S_{xx}=\sum_{1}^{7}x^2_i-n\bar{x}^2

= 224 - 7*4 *4

= 112

100(1 - \alpha )% confidence interval for  \beta_1 is

\beta_1\pm t_{\alpha/2}n-2\sqrt{\frac{MS_{Res}}{S_{xx}}}

i.e,

P(\hat{\beta_1}-t_{\alpha/2}*n-2\sqrt{\frac{MS_{Residual}}{S_{xx}}}\leq \beta_1 \leq \hat{\beta_1}+t_{\alpha/2}*n-2\sqrt{\frac{MS_{Residual}}{S_{xx}}})=1-\alpha

\Rightarrow P(0.9464-2.571\sqrt{\frac{2.2213}{112}}\leq \beta_1 \leq 0.9464+2.571\sqrt{\frac{2.2213}{112}})=0.95

\Rightarrow P(0.9464-2.571*0.1408\leq \beta_1 \leq 0.9464+2.571*0.1408)=0.95 \, \, \, \, \, \left \{ \because t_{0.05/2},7-2=2.571\, \, \alpha=0.05 \right \}

0.9464 \pm 0.3619968

0.9464 \pm 0.36200

Here I take  \hat{\beta_1}=0.9464, Take  \hat{\beta_1}=0.94643 and perfrom calculation then you will get exact 0.9463 \pm 0.36203 95% confidence interval.

The correct answer is option (D)

Thank you for your supporting. Please upvote my answer...